OpenStudy (anonymous):

which line is perpendicular to the line y=3/4x+11

OpenStudy (anonymous):

option..??

OpenStudy (anonymous):

but keep in mind any two lines are perpndicular if product of thee slopes = -1 to find slope compare each of your eq in choic and in question with this. y=mx+c here m=slope

OpenStudy (anonymous):

and slope os that line may be = -4/3

OpenStudy (anonymous):

y=3/4x-11, -3/4x+11, 4/3x+9, or -4/3x+9 those are my options to choose from'

OpenStudy (anonymous):

so optin B is right choice...see the slope=-3/4 the only choice aith this...value...

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome...